OpenStudy (zyberg):

Determine the sign of c, if equality ax^2 + bx + c = 0 doesn't have any solutions and inequality a + b + c < 0 is right.

OpenStudy (hawk):

Is it ax saquare?

OpenStudy (hawk):

square

OpenStudy (zyberg):

Yes. \[ax^2 + bx + c = 0; a+ b + c < 0\]

OpenStudy (hawk):

you want sign of c right?

OpenStudy (zyberg):

I understand that from ax^2 + bx + c = 0 not having any solutions, I can write that: b^2 - 4ac < 0 But I don't see how I would obtain the sign of c in this way.

OpenStudy (zyberg):

@Hawk, yes.

OpenStudy (zyberg):

ax^2 + bx + c = 0 Doesn't have solutions only when a and c have same signs.

OpenStudy (hawk):

Yep when we take \[b^{2}-4ac\]

OpenStudy (hawk):

Sign of c can be positive i think

OpenStudy (zyberg):

Just noticed that, if a and c are negative, then inequality b < a + c holds true.

OpenStudy (hawk):

Yep

OpenStudy (hawk):

So ax

OpenStudy (hawk):

\[a^{x}2 +b+c=0\]

OpenStudy (zyberg):

Sorry, but why?

ganeshie8 (ganeshie8):

.

OpenStudy (zyberg):

@ganeshie8 Hello! Do you have any ideas?

ganeshie8 (ganeshie8):

I might have some idea, but I'm busy in kitchen atm. Will give this a good try after dinner :)

OpenStudy (zyberg):

@ganeshie8 thank you very much! :)

OpenStudy (mayankdevnani):

well `c` must be less than zero

OpenStudy (welshfella):

4ac > b^2 so either a and c are both negative or they are both positive

OpenStudy (mayankdevnani):

 for this equation, a<0 x co-ordinate of vertex <0 \[\large \bf \frac{-b}{2a}<0\] If a<0,then b should be less than zero to satisfy it i.e b<0 To prove this a + b + c < 0 For good solution, c must be less than zero

OpenStudy (mayankdevnani):

and you can see that y-intercept of the graph is negative ! therefore, c<0

OpenStudy (mayankdevnani):

what do you think? @welshfella

OpenStudy (welshfella):

looks good

Parth (parthkohli):

As an alternative, observe that \(a+b+c= p(1)\). Now since your curve must maintain the same sign (otherwise, if it doesn't, then it will have a root by IVT) we can say that \(p(0)\) has the same sign as \(p(1)\). But \(p(0)= c\)...

ganeshie8 (ganeshie8):

IVT worked here like a charm !